### v1

@@ 5.4 Comparison with State-of-the-art Methods
Results on DSText. We further conducted experiments on DSText, focusing on dense and small video text scenarios, with results shown in Tab. 7. Notably, many top methods on the DSText leaderboard rely on model ensembles and large public datasets to boost performance [34]. For instance, TencentOCR combines detection outputs from DBNet [24] and Cascade MaskRCNN [7], integrates the Parseq [2] text recognizer, and improves end-to-end tracking with ByteTrack [40]. Similarly, DA uses Mask R-CNN [11] and DBNet for text detection, BotSORT [1] to replace the tracker in VideoTextSCM [10], and Parseq for recognition. As a single model, our GoMatching++ slightly trails behind ensemble-based methods in video text tracking. However, this performance gap can be reduced with stronger detectors. Our method shows strong video text spotting capabilities, ranking first in MOTA (23.23%), second in MOTP (80.43%), and third in IDF1 (46.24%) for video text spotting. Compared to the best single-model approach, GoMatching++ achieves significant improvements: +9.04% in MOTA and +2.83% in IDF1 for video text tracking, and +45.86% in MOTA and +19.81% in IDF1 for video text spotting.
More Comparisons between TransDETR and GoMatching++. We further present additional comparisons between our methods and TransDETR on ICDAR15-video. In the first row of Tab. 8, the input image’s shorter side is set to 800, matching the default configuration of TransDETR. As shown in Tab. 8, GoMatching consistently outperforms TransDETR across all settings. Notably, it exceeds TransDETR in both inference speed and spotting accuracy under its default configuration. When the image size increases (e.g., to 1,000), GoMatching offers superior spotting performance, albeit with a slight reduction in inference speed. With the lighter LST-Matcher design, GoMatching++ achieves better performance and faster inference, highlighting its efficiency and effectiveness. Additionally, we compare the number of trainable parameters and training costs in Tab. 9. GoMatching++ requires significantly less training budget than TransDETR due to its simpler architecture, with a 70% reduction in trainable parameters, 19.4GB less memory usage, and 301 fewer GPU-hours. This makes GoMatching++ a more efficient and competitive baseline for video text spotting.
5.5 Failure Cases Analysis
@@
Table 10: Intra-dataset evaluation results. Models are trained on the ArTVideo training set and evaluated on the ArTVideo test set. $\bigstar$’ indicates that the structure of TransDETR is modified to utilize polygon and mask annotations in ArTVideo.
@@
TransDETR-mask (ArTVideo) 65.5(+11.3) 82.3(+14.4) 77.6(+7.2) 21.1(+18.3) 86.2(+16.8) 59.3(+10.0)
@@
Table 11: Inter-dataset evaluation results. ‘*’: Model evaluated using the officially released version. $\bigstar$’: TransDETR structure modified to support polygon and mask annotations in ArTVideo.
6 Experiments on ArTVideo
6.1 Implementation Details
We train GoMatching and GoMatching++ under the same settings outlined in Sec. 5.2, using a 1280-pixel input image size during inference. The original TransDETR model, designed for quadrilateral bounding boxes, does not fully leverage the fine-grained annotations (polygons and masks) in ArTVideo. To address this, we enhance TransDETR with an additional mask branch, inspired by Mask R-CNN [11], creating a new variant called TransDETR-mask. This modification enables the model to utilize mask annotations from ArTVideo, improving the accuracy of text contour predictions. Apart from this change, all other settings follow the default configuration in the open-source implementation.
6.2 Intra-Dataset Evaluation
To evaluate the effectiveness of our methods for curved text, we compare GoMatching and GoMatching++ with TransDETR-mask on the ArTVideo dataset. As shown in Tab. 10, both GoMatching and GoMatching++ outperform TransDETR-mask across all evaluation tasks. The performance gap is particularly notable when adding a recognition task for end-to-end spotting or focusing solely on curved text. GoMatching++ achieves a 98.8% improvement in MOTA and a 33.0% improvement in IDF1 over TransDETR-mask. These results highlight the limitations of prior SOTA methods, which struggle with recognition and adaptability due to the inherent design of their frameworks. Furthermore, the performance boost of GoMatching++ over GoMatching emphasizes the effectiveness of its optimized design. Visualization results in Fig. 10 demonstrate a clear improvement in text recognition by GoMatching++ compared to TransDETR-mask.
6.3 Inter-Datasets Evaluation
In this subsection, we explore the transferability and generalization of models across datasets through inter-dataset evaluations between the ICDAR15-video and ArTVideo datasets. As shown in Tab. 11, the lack of curved text in ICDAR15-video results in significantly lower performance when models trained on ICDAR15-video are tested on ArTVideo, especially for tasks involving curved text. For example, TransDETR, trained on ICDAR15-video, experiences a 37.7% drop in MOTA and 21.4% in IDF1 for video curved text tracking, and a 41.2% drop in MOTA and 20.1% in IDF1 for video curved text spotting compared to models trained on ArTVideo. GoMatching and GoMatching++ perform better, although they still experience a 10% drop in MOTA due to the lack of curved text in the training data. In contrast, models trained on ArTVideo and tested on ICDAR15-video perform similarly to those trained directly on ICDAR15-video, with performance decreases not exceeding 4%, and some metrics even showing improvements. These results demonstrate GoMatching++’s superiority over prior SOTA methods and highlight the diversity of the proposed ArTVideo dataset and its potential to enhance model robustness across various video text spotting tasks.
7 Conclusion
@@
[…]
Acknowledgements This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFC2705700, in part by the National Natural Science Foundation of China under Grants U23B2048 and 62225113, in part by the Innovative Research Group Project of Hubei Province under Grant 2024AFA017, and in part by the Science and Technology Major Project of Hubei Province under Grant 2024BAB046. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
Declarations
  • Funding. Dr. Juhua Liu was supported by the National Natural Science Foundation of China under Grants U23B2048 and 62076186, and the Science and Technology Major Project of Hubei Province under Grant 2024BAB046.
    Dr. Bo Du was supported by the National Key Research and Development Program of China under Grant 2023YFC2705700, the National Natural Science Foundation of China under Grants 62225113, and the Innovative Research Group Project of Hubei Province under Grant 2024AFA017.
  • Competing Interests. The authors have no relevant financial or non-financial interests to disclose.
  • Data Availability. All datasets used in this paper are publicly available. ICDAR13-video [16] and ICDAR15-video [14] can be accessed at https://rrc.cvc.uab.es/?ch=3. DSText [34] can be accessed at https://rrc.cvc.uab.es/?ch=22. BOVText [33] can be accessed at https://github.com/weijiawu/BOVText-Benchmark. ArTVideo can be accessed at https://github.com/Hxyz-123/GoMatching.

### v2

@@ V-D Comparison with State-of-the-art Methods
More Comparisons between TransDETR and GoMatching++. We further present additional comparisons between our methods and TransDETR on ICDAR15-video. In the first row of Tab. VII, the input image’s shorter side is set to 800, matching the default configuration of TransDETR. As shown in Tab. VII, GoMatching consistently outperforms TransDETR across all settings. Notably, it exceeds TransDETR in both inference speed and spotting accuracy under its default configuration. When the image size increases (e.g., to 1,000), GoMatching offers superior spotting performance, albeit with a slight reduction in inference speed. With the lighter LST-Matcher design, GoMatching++ achieves better performance and faster inference, highlighting its efficiency and effectiveness. Additionally, we compare the number of trainable parameters and training costs in Tab. VII. GoMatching++ requires significantly less training budget than TransDETR due to its simpler architecture, with a 70% reduction in trainable parameters, 19.4GB less memory usage, and 301 fewer GPU-hours. This makes GoMatching++ a more efficient and competitive baseline for video text spotting.
V-E Failure Cases Analysis
@@
TABLE VIII: Intra-dataset evaluation results. Models are trained on the ArTVideo training set and evaluated on the ArTVideo test set. $\bigstar$’ indicates that the structure of TransDETR is modified to utilize polygon and mask annotations in ArTVideo.
VI Experiments on ArTVideo
VI-A Implementation Details
We train GoMatching and GoMatching++ under the same settings outlined in Sec. V-B, using a 1280-pixel input image size during inference. The original TransDETR model, designed for quadrilateral bounding boxes, does not fully leverage the fine-grained annotations (polygons and masks) in ArTVideo. To address this, we enhance TransDETR with an additional mask branch, inspired by Mask R-CNN [he2017mask], creating a new variant called TransDETR-mask. This modification enables the model to utilize mask annotations from ArTVideo, improving the accuracy of text contour predictions. Apart from this change, all other settings follow the default configuration in the open-source implementation.
@@ VI-A Implementation Details
TransDETR-mask (ArTVideo) 65.5(+11.3) 82.3(+14.4) 74.6(+4.2) 21.1(+18.3) 86.2(+16.8) 59.3(+10.0)
@@
TABLE IX: Inter-dataset evaluation results. ‘*’: Model evaluated using the officially released version. $\bigstar$’: TransDETR structure modified to support polygon and mask annotations in ArTVideo.
VI-B Intra-Dataset Evaluation
To evaluate the effectiveness of our methods for curved text, we compare GoMatching and GoMatching++ with TransDETR-mask on the ArTVideo dataset. As shown in Tab. VIII, both GoMatching and GoMatching++ outperform TransDETR-mask across all evaluation tasks in curved-text setting. The performance gap is particularly notable when adding a recognition task for end-to-end spotting or focusing solely on curved text. GoMatching++ achieves a 98.8% improvement in MOTA and a 33.0% improvement in IDF1 over TransDETR-mask. These results highlight the limitations of prior SOTA methods, which struggle with recognition and adaptability due to the inherent design of their frameworks. Furthermore, the performance boost of GoMatching++ over GoMatching emphasizes the effectiveness of its optimized design. Visualization results in Fig. 9 demonstrate a clear improvement in text recognition by GoMatching++ compared to TransDETR-mask.
VI-C Inter-Datasets Evaluation
In this subsection, we explore the transferability and generalization of models across datasets through inter-dataset evaluations between the ICDAR15-video and ArTVideo datasets. As shown in Tab. IX, the lack of curved text in ICDAR15-video results in significantly lower performance when models trained on ICDAR15-video are tested on ArTVideo, especially for tasks involving curved text. For example, TransDETR, trained on ICDAR15-video, experiences a 37.7% drop in MOTA and 21.4% in IDF1 for video curved text tracking, and a 41.2% drop in MOTA and 20.1% in IDF1 for video curved text spotting compared to models trained on ArTVideo. GoMatching and GoMatching++ perform better, although they still experience a 10% drop in MOTA due to the lack of curved text in the training data. In contrast, models trained on ArTVideo and tested on ICDAR15-video perform similarly to those trained directly on ICDAR15-video, with performance decreases not exceeding 4%, and some metrics even showing improvements. These results demonstrate GoMatching++’s superiority over prior SOTA methods and highlight the diversity of the proposed ArTVideo dataset and its potential to enhance model robustness across various video text spotting tasks.
VII Conclusion
@@
Acknowledgements
This work was supported by WHU-Kingsoft Joint Lab. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.